# GENPACK: KPI-Guided Multi-Objective Genetic Algorithm for Industrial 3D Bin Packing

### v2

@@ 4.5. Parameters
6: $S_{GA}\leftarrow\arg\max_{C\in\mathbf{P}_{0}}f(C)$; ; $\mathbf{P}\leftarrow\mathbf{P}_{0}$
@@ 5. Hybrid Pipeline Architecture
This section outlines the overall workflow that combines constructive heuristics, genetic refinement, and post-processing into a unified pipeline. Our method follows a two-phase placement strategy: in the first phase, constructive heuristics generate an initial packing using MaxRects, while the second phase applies a genetic algorithm to refine the residual items missed by MaxRects through KPI-weighted optimization before merging them back into the final configuration.
@@ 5.1. Superitem Preprocessing
To reduce the effective search space, similar items are grouped into superitems for the heuristics (elhedhli2019three, ). Three categories are considered: Single superitems represent individual items. Horizontal superitems combine multiple items of identical height. Vertical superitems stack compatible items into stable columns up to a maximum height. These groupings reflect practical packing patterns and improve scalability by reducing the number of decision variables while preserving feasible packing opportunities.
@@ 5.2. Integration Strategy
After GA optimization, residual packings are merged with the constructive baseline (Section 4). Layers are aligned with the stacking height produced in the first phase, and all placements are validated to ensure feasibility (Section 3.2), KPI scores are then recomputed for the combined solution.
@@ 5.3. Post-Processing
Although GA refinement improves density, stability, and balance, it does not always ensure full feasibility, as defined in Section 3.2. A small fraction of items may remain unplaced, or minor density irregularities may weaken overall packing quality. Post-processing addresses these issues through a multi-phase heuristic procedure (Algorithm 2). The mathematical formulation for the post-processing is discussed in detail in Appendix A.
The procedure begins with directional compaction, where items in each layer are iteratively shifted toward the pallet origin to minimize empty space while preserving support stability. In the second phase, fallback placement handles any unplaced items by evaluating feasible positions near pallet boundaries and, if necessary, performing a coarse grid search with rotation to recover remaining items. Finally, a validation stage ensures full feasibility by removing overlapping items or those failing the minimum support threshold.
@@ 5.4. Computational Complexity
Per generation, we therefore spend $\Theta(P\,n^{2})$ on fitness, plus $O(P\log P)$ for elite selection and up to $O(P^{2})$ from tournament selection in reproduction. For practical settings where $n$ is moderate to large, the overall cost scales as:
@@ 6.3. Baselines
We compared GENPACK against seven algorithms covering classical heuristics, learning-based methods, and our constructive–hybrid variants (Table 1). External baselines include O3D-BPP-PCT (zhao2022pct, ), GOPT (xiong2024gopt, ), Extreme Point (crainic2008extreme, ), and Sisyphus (demisse2012mixed, ). Internal baselines include MaxRects (jylanki2010maxrects, ), Hybrid-GA, and Hybrid-GA (+PP), which represent successive stages of our pipeline and isolate the effect of each component.
@@
All experiments were conducted on a machine with Intel Core Ultra 7 155U processor (1.70 GHz) with 32 GB RAM running Windows 11 Pro. The full code implementation will be made available for reproducibility soon.
@@
GOPT (xiong2024gopt, ) Learning-based Deep Reinforcement learning method using Transformer architecture.
@@
Hybrid-GA (core) Hybrid (ours, ablation) Extends MaxRects with GA refinement of residuals but without post-processing.
Hybrid-GA (+PP) Hybrid (ours, final) Full proposed pipeline: MaxRects, GA refinement, and post-processing.
@@
Hybrid-GA (core) 0.538 $\pm$ 0.082 0.657 $\pm$ 0.072 0.696 $\pm$ 0.049
Hybrid-GA (+PP) 0.469 $\pm$ 0.098 0.531 $\pm$ 0.119 0.651 $\pm$ 0.087
@@
Hybrid-GA (core) 0.561 $\pm$ 0.076 0.583 $\pm$ 0.121 0.851 $\pm$ 0.057 0.748 $\pm$ 0.071
Hybrid-GA (+PP) 0.382 $\pm$ 0.099 0.815 $\pm$ 0.114 0.803 $\pm$ 0.111 0.670 $\pm$ 0.110
@@
Hybrid-GA (core) 3.84 2.15
Hybrid-GA (+PP) 29.20 28.09
@@ 7.1. KPI-Based Analysis
Table 2 shows that Hybrid-GA (core) achieves the highest scores in absolute and relative density as well as height–width ratio, confirming that GA refinement effectively compacts items and reduces vertical instability. It also exhibits the lowest standard deviations across all efficiency metrics, demonstrating superior consistency and robustness across diverse order configurations. In contrast, pure heuristics such as MaxRects underperform in density, while Extreme Point achieves moderately good stability but lags in utilization. The DRL baselines (O3D-BPP-PCT, GOPT) show high variance, indicating less predictable performance.
@@
As shown in Table 4, GA refinement also strengthens balance, with Hybrid-GA (core) achieving the best overall center of gravity alignment. Hybrid-GA (+PP) attains the highest surface support, reflecting its ability to improve contact stability across layers, albeit at the cost of reduced side support, indicating a trade-off between lateral and vertical stability. Beyond mean performance, Hybrid-GA (core) demonstrates the lowest variability in center of gravity measures and strong consistency, confirming that GA-based optimization yields reliable and stable balance across different problem instances. Learning-based methods (O3D-BPP-PCT, GOPT) achieve moderate performance but remain consistently weaker than GA-based pipelines, exhibiting on average larger variance, suggesting less predictable solution quality.
@@
In addition to aggregate KPI scores, Figure 2(b) shows the evolution of absolute density and surface support as a function of the number of successfully packed items. The histogram (bottom) reveals that order sizes are concentrated in the range of 15–60 items, with the distribution peaking in the 25–35 item range and maintaining substantial representation through 45 items, thus motivating our focused analysis. On the efficiency side (top), O3D-BPP-PCT starts from low density and shows only limited improvement across the range, while GOPT and Sisyphus remain consistently and significantly below densities of Hybrid-GA (+PP) and Extreme point. The latter attains comparable density trends but remains below Hybrid-GA (+PP) on average, which maintains steady improvement and achieves the highest density across this critical range. On the stability side (middle), O3D-BPP-PCT remains relatively weak throughout, GOPT and Sisyphus similarly underperform, and Extreme Point, though more stable, achieves lower surface support compared to Hybrid-GA (+PP). The latter exhibits both higher and flatter curves, reflecting its ability to preserve surface contact and structural robustness across varying order sizes. These qualitative trends supplement the tabular results in Table 4 by showing that hybrid GA pipelines not only achieve stronger KPI averages but also maintain robustness across the distribution of order sizes.
@@
Figure 2(a) shows representative packings. Learning-based baselines (O3D-BPP-PCT, GOPT) often produce tall and irregular stacks with noticeable gaps on this real-world industrial dataset, while heuristic methods (Sisyphus, Extreme Point) yield more compact structures but with vertical instability or uneven support. While Sisyphus attains surface support values comparable to those of Hybrid-GA (+PP), it frequently produces tall, unstable stacks, proving that numerical support alone does not guarantee feasible industrial packings. In contrast, the hybrid GA pipelines generate more balanced and layered configurations, with Hybrid-GA (+PP) achieving higher compaction and stronger surface support.
@@
Inference time is crucial for large-scale industrial use, where thousands of orders are processed daily. Table 4 reports statistics for execution times across 1500 orders. While heuristic and learning-based methods achieved sub-second to few-second runtimes, Hybrid-GA variants required longer execution (~4–29 s per order) due to their multi-stage optimization.
@@ 8. Discussion
While heuristics and learning-based approaches process orders nearly instantaneously but at the expense of solution quality. In contrast, the hybrid GA (+PP) requires longer runtimes up to 29 seconds per order for the full pipeline. In warehouse contexts, where optimization can be run offline or in parallel, these runtimes can be considered acceptable given the significant improvements in packing feasibility.
Overall, the results indicate that the primary bottleneck in industrial deployment is not speed but solution quality under operational constraints. Fast heuristics and neural methods can deliver throughput but often fail to guarantee stability or balanced load distribution. By explicitly optimizing multiple KPIs, our method achieves configurations that are both efficient and safe, even if this is achieved at higher computational cost.
@@ 9. Conclusion and Future Work
We presented a genetic algorithm (GA) tailored to industrial 3D bin packing and embedded it within a hybrid pipeline combining constructive heuristics and post-processing. The method employs a KPI-driven fitness function addressing utilization, stability, support, and balance; a layer-based chromosome preserving feasibility; and structure-aware operators. These components enable the GA to perform global refinements beyond rule-based placement while remaining compatible with palletization constraints.
Compared with constructive strategies such as MaxRects (jylanki2010maxrects, ) and Extreme Point (crainic2008extreme, ), which prioritize speed and initial utilization, the GA provides incremental improvements in balance, stability, and support. Its role is not to generate full packings but to refine them toward industrially relevant criteria. These improvements come with slightly higher computational cost, making the overall benefit dependent on latency budgets and the chosen KPI weighting.
@@ Appendix B Additional KPI trends
Figure 3 (top) compares Relative Density and Height–Width Ratio across methods. Hybrid-GA (+PP) achieves consistently high relative density and remains at or above the baselines across most of the range of packed items. O3D-BPP-PCT starts lower but steadily improves, approaching Hybrid-GA (+PP) at higher item counts; Extreme-Point remains mid-range, while Sisyphus and GOPT fluctuate and dip more frequently. For Height–Width Ratio, Hybrid-GA (+PP) and Extreme-Point maintain comparatively higher and more stable ratios, whereas O3D-BPP-PCT, Sisyphus, and GOPT exhibit greater variance, indicating less consistent vertical compactness.
@@
In the lower plots of Figure 3, Side Support generally increases with the number of packed items. Hybrid-GA (+PP) maintains consistent support throughout, whereas O3D-BPP-PCT struggles in the lower-item range but recovers to achieve the strongest support at higher item counts. Extreme-Point follows closely behind, while GOPT remains lowest on average. CoG2D values for all methods cluster near the center as item count increases; Hybrid-GA (+PP) exhibits slightly tighter dispersion, indicating steadier horizontal balance. CoG3D shows a similar convergence trend at higher counts, with early-bin variability more pronounced for Sisyphus and GOPT. Overall, Hybrid-GA (+PP) is competitive or leading in both compactness and stability, while O3D-BPP-PCT improves markedly as instance size grows.

### v3

@@ 4.5. Parameters
6: $S_{GA}\leftarrow\arg\max_{C\in\mathbf{P}_{0}}f(C)$; $\mathbf{P}\leftarrow\mathbf{P}_{0}$
@@ 5. Hybrid Pipeline Architecture
This section outlines the overall workflow that combines constructive heuristics, genetic refinement, and post-processing into a unified pipeline. Our method follows a three-stage placement strategy: in the first stage, constructive heuristics generate an initial packing using MaxRects; in the second stage, a genetic algorithm refines the residual items missed by MaxRects through KPI-weighted optimization; and in the final stage, the refined solution is merged back into the full configuration and post-processed.
@@ 5.1. Superitem Preprocessing
To reduce the effective search space, similar items are grouped into superitems for the heuristics (elhedhli2019three, ). Three categories are considered: single superitems represent individual items, horizontal superitems combine multiple items of identical height, and vertical superitems stack compatible items into stable columns up to a maximum height. These groupings reflect practical packing patterns and improve scalability by reducing the number of decision variables while preserving feasible packing opportunities.
@@ 5.2. Integration Strategy
After GA optimization, residual packings are merged with the constructive baseline (Section 4). Layers are aligned with the stacking height produced in the first stage, and all placements are validated to ensure feasibility (Section 3.2); KPI scores are then recomputed for the combined solution.
@@ 5.3. Post-Processing
Although GA refinement improves density, stability, and balance, it does not always ensure full feasibility, as defined in Section 3.2. A small fraction of items may remain unplaced, or minor density irregularities may weaken overall packing quality. Post-processing addresses these issues through a multi-stage heuristic procedure (Algorithm 2). The mathematical formulation of the post-processing is discussed in detail in Appendix A.
The procedure begins with directional compaction, where items in each layer are iteratively shifted toward the pallet origin to minimize empty space while preserving support stability. In the second stage, fallback placement handles any unplaced items by evaluating feasible positions near pallet boundaries and, if necessary, performing a coarse grid search with rotation to recover remaining items. Finally, a validation stage ensures full feasibility by removing overlapping items or those failing the minimum support threshold.
@@ 5.4. Computational Complexity
Per generation, we therefore spend $\Theta(P\,n^{2})$ on fitness, plus $O(P\log P)$ for elite selection and up to $O(P^{2})$ from tournament selection in reproduction. For practical settings where $n$ is moderate to large, the overall cost scales as
@@ 6.3. Baselines
We compared GENPACK against seven algorithms covering classical heuristics, learning-based methods, and our constructive–hybrid variants (Table 1). External baselines include O3D-BPP-PCT (zhao2022pct, ), GOPT (xiong2024gopt, ), Extreme Point (crainic2008extreme, ), and Sisyphus (demisse2012mixed, ). Internal baselines include MaxRects (jylanki2010maxrects, ), GENPACK, and GENPACK (+PP), which represent successive stages of our pipeline and isolate the effect of each component.
@@
All experiments were conducted on a machine with an Intel Core Ultra 7 155U processor (1.70 GHz), 32 GB RAM, and Windows 11 Pro. The full implementation is available at GitHub for reproducibility.
@@
GOPT (xiong2024gopt, ) Learning-based Deep reinforcement learning method using a Transformer architecture.
@@
GENPACK (core) Hybrid (ours, ablation) Extends MaxRects with GA refinement of residuals but without post-processing.
GENPACK (+PP) Hybrid (ours, final) Full proposed pipeline: MaxRects, GA refinement, and post-processing.
@@
GENPACK (core) 0.538 $\pm$ 0.082 0.657 $\pm$ 0.072 0.696 $\pm$ 0.049
GENPACK (+PP) 0.469 $\pm$ 0.098 0.531 $\pm$ 0.119 0.651 $\pm$ 0.087
@@
GENPACK (core) 0.561 $\pm$ 0.076 0.583 $\pm$ 0.121 0.851 $\pm$ 0.057 0.748 $\pm$ 0.071
GENPACK (+PP) 0.382 $\pm$ 0.099 0.815 $\pm$ 0.114 0.803 $\pm$ 0.111 0.670 $\pm$ 0.110
@@
GENPACK (core) 3.84 2.15
GENPACK (+PP) 29.20 28.09
@@ 7.1. KPI-Based Analysis
Table 2 shows that GENPACK (core) achieves the highest scores in absolute and relative density as well as height–width ratio, confirming that GA refinement effectively compacts items and reduces vertical instability. It also exhibits the lowest standard deviations across all efficiency metrics, demonstrating superior consistency and robustness across diverse order configurations. In contrast, pure heuristics such as MaxRects underperform in density, while Extreme Point achieves moderately good stability but lags in utilization. The DRL baselines (O3D-BPP-PCT, GOPT) show high variance, indicating less predictable performance.
@@
As shown in Table 4, GA refinement also strengthens balance, with GENPACK (core) achieving the best overall center-of-gravity alignment. GENPACK (+PP) attains the highest surface support, reflecting its ability to improve contact stability across layers, albeit at the cost of reduced side support, indicating a trade-off between lateral and vertical stability. Beyond mean performance, GENPACK (core) demonstrates the lowest variability in center-of-gravity measures and strong consistency, confirming that GA-based optimization yields reliable and stable balance across different problem instances. Learning-based methods (O3D-BPP-PCT, GOPT) achieve moderate performance but remain consistently weaker than GA-based pipelines, exhibiting on average larger variance and thus less predictable solution quality.
@@
In addition to aggregate KPI scores, Figure 2(b) shows the evolution of absolute density and surface support as a function of the number of successfully packed items. The histogram (bottom) reveals that order sizes are concentrated in the range of 15–60 items, with the distribution peaking in the 25–35 item range and maintaining substantial representation through 45 items, thus motivating our focused analysis. On the efficiency side (top), O3D-BPP-PCT starts from low density and shows only limited improvement across the range, while GOPT and Sisyphus remain consistently and significantly below the densities of GENPACK (+PP) and Extreme Point. The latter attains comparable density trends but remains below GENPACK (+PP) on average, which maintains steady improvement and achieves the highest density across this critical range. On the stability side (middle), O3D-BPP-PCT remains relatively weak throughout, GOPT and Sisyphus similarly underperform, and Extreme Point, though more stable, achieves lower surface support than GENPACK (+PP). The latter exhibits both higher and flatter curves, reflecting its ability to preserve surface contact and structural robustness across varying order sizes. These qualitative trends supplement the tabular results in Table 4 by showing that GENPACK pipelines not only achieve stronger KPI averages but also maintain robustness across the distribution of order sizes.
@@
Figure 2(a) shows representative packings. Learning-based baselines (O3D-BPP-PCT, GOPT) often produce tall and irregular stacks with noticeable gaps on this real-world industrial dataset, while heuristic methods (Sisyphus, Extreme Point) yield more compact structures but with vertical instability or uneven support. While Sisyphus attains surface-support values comparable to those of GENPACK (+PP), it frequently produces tall, unstable stacks, showing that numerical support alone does not guarantee feasible industrial packings. In contrast, the GENPACK pipelines generate more balanced and layered configurations, with GENPACK (+PP) achieving higher compaction and stronger surface support.
@@
Inference time is crucial for large-scale industrial use, where thousands of orders are processed daily. Table 4 reports statistics for execution times across 1500 orders. While heuristic and learning-based methods achieved sub-second to few-second runtimes, GENPACK variants required longer execution (~4–29 s per order) due to their multi-stage optimization.
@@ 8. Discussion
While heuristics and learning-based approaches process orders nearly instantaneously, they do so at the expense of solution quality. In contrast, the GENPACK (+PP) requires longer runtimes, up to 29 seconds per order for the full pipeline. In warehouse contexts, where optimization can be run offline or in parallel, these runtimes can still be considered acceptable given the significant improvements in packing feasibility.
Overall, the results indicate that the primary bottleneck in industrial deployment is not speed alone, but solution quality under operational constraints. Fast heuristics and neural methods can deliver throughput but often fail to guarantee stability or balanced load distribution. By explicitly optimizing multiple KPIs, our method achieves configurations that are both efficient and safe, even if this comes at higher computational cost.
@@ 9. Conclusion and Future Work
We presented a genetic algorithm (GA) tailored to industrial 3D bin packing and embedded it within a hybrid pipeline combining constructive heuristics and post-processing. The method employs a KPI-guided fitness function addressing utilization, stability, support, and balance; a layer-based chromosome preserving feasibility; and structure-aware operators. These components enable the GA to perform global refinements beyond rule-based placement while remaining compatible with palletization constraints.
Compared with constructive strategies such as MaxRects (jylanki2010maxrects, ) and Extreme Point (crainic2008extreme, ), which prioritize speed and initial utilization, the GA provides incremental improvements in balance, stability, and support. Its role is not to generate full packings, but to refine them toward industrially relevant criteria. These improvements come with slightly higher computational cost, making the overall benefit dependent on latency budgets and the chosen KPI weighting.
@@ Appendix B Additional KPI trends
Figure 3 (top) compares Relative Density and Height–Width Ratio across methods. GENPACK (+PP) achieves consistently high relative density and remains at or above the baselines across most of the range of packed items. O3D-BPP-PCT starts lower but steadily improves, approaching GENPACK (+PP) at higher item counts; Extreme Point remains mid-range, while Sisyphus and GOPT fluctuate and dip more frequently. For Height–Width Ratio, GENPACK (+PP) and Extreme Point maintain comparatively higher and more stable ratios, whereas O3D-BPP-PCT, Sisyphus, and GOPT exhibit greater variance, indicating less consistent vertical compactness.
@@
In the lower plots of Figure 3, Side Support generally increases with the number of packed items. GENPACK (+PP) maintains consistent support throughout, whereas O3D-BPP-PCT struggles in the lower-item range but recovers to achieve the strongest support at higher item counts. Extreme Point follows closely behind, while GOPT remains lowest on average. CoG2D values for all methods cluster near the center as item count increases; GENPACK (+PP) exhibits slightly tighter dispersion, indicating steadier horizontal balance. CoG3D shows a similar convergence trend at higher counts, with early-bin variability more pronounced for Sisyphus and GOPT. Overall, GENPACK (+PP) is competitive or leading in both compactness and stability, while O3D-BPP-PCT improves markedly as instance size grows.